find the equation of the parabola whose vertex is at (h,k) and satisfying the given conditions. opens upward, length of the latus rectum =8, and passing through (1, 1/2) and (7, -1).
Answer by lwsshak3(11628)

find the equation of the parabola whose vertex is at (h,k) and satisfying the given conditions. opens upward, length of the latus rectum =8, and passing through (1, 1/2) and (7, -1).
***
basic form of equation for a parabola: (x-h)^2=4p(y-k)
latus rectum=8=4p
equation: (x-h)^2=8(y-k)
2-equations using coordinates of 2 given points (1, 1/2) and (7, -1).
..
(1-h)^2=8(1/2-k)
(7-h)^2=8(-1-k)
..
1-2h+h^2=4-8k
49-14h+h^2=-8-8k
subtract eliminating h^2 and-8k
-48+12h=12
12h=60
h=5
8k=4-1+2h-h^2=3+10-25=-12
k=-12/8=-3/2=-1.5
equation: (x-5)^2=8(y+1.5)